Clayton is 2 years younger than his brother Josh. If the sum of their ages is 20, how old is each boy?
Found 2 solutions by algebrahouse.com, stanbon:
Answer by algebrahouse.com(1659)

j = Josh's age
j - 2 = Clayton's age {Clayton is 2 years younger than Josh}

j + j - 2 = 20 {sum of their ages is 20}
2j - 2 = 20 {combined like terms}
2j = 22 {added 2 to each side}
j = 11 {divided each side by 2}
j - 2 = 9 {substituted 11, in for j, into j - 2}

Josh is 11
Clayton is 9

For more help from me, visit: www.algebrahouse.com



Answer by stanbon(75887)

Clayton is 2 years younger than his brother Josh. If the sum of their ages is 20, how old is each boy?
-----
c = j-2
c + j = 20
----------
Substitute for "c" and solve for "j"::
j-2 + j = 20
2j = 22
josh = 11 yrs
------
clayton = j-2 = 9 yrs
------------------
Cheers,
Stan H.
------------